OpenStudy (anonymous):

Use the dot product to find the magnitude of u if u = 6i - 3j

OpenStudy (nubeer):

\[magnitude = \sqrt{6^{2}+(-3^{2})}\]

OpenStudy (anonymous):

I solved that and I got sqrt of 45..which is not one of my answers

OpenStudy (anonymous):

i mean its not one of my choices

OpenStudy (nubeer):

what are your choices?

OpenStudy (anonymous):

18, 4sqrt 5, 3, 23, 3sqrt 5

OpenStudy (anonymous):

√45 = 3√5

OpenStudy (nubeer):

\[\sqrt{45} =\sqrt{9\times5}=\sqrt{9}\times \sqrt{5}=3\sqrt{5}\]

OpenStudy (anonymous):

thank you!